The weight of competition pumpkins at the Circleville Pumpkin Show in Circleville, Ohio, can be represented by a normal distribution with a mean of 703 pounds and a standard deviation of 347 pounds.

a. Find the probability that a randomly selected pumpkin weighs at least 1,622 pounds.

b. Find the probability that a randomly selected pumpkin weighs between 465.1 and 1,622

pounds.


1
Expert's answer
2020-07-09T18:22:04-0400

Let r. v. "X" be the weight of randomly choosen pumpkin. Then "X\in N (703;347^2)."

a. We should find "P\{X\geq 1622\}."

"P\{X\geq 1622\}=1-P\{X<1622\}."

"P\{X<1622\}=F(1622)" where "F(x)=\frac{1}{\sqrt{2\pi}347}\int_{-\infty}^x e^{-\frac{(t-703)^2}{2\cdot 347^2}}dt."

"F(1622)=\Phi(\frac{1622-703}{347})\approx\Phi(2.648)\approx 0.9960" where

"\Phi(x)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^x e^{-\frac{z^2}{2}}dz."

So "P\{X\geq 1622\}=1-P\{X<1622\}\approx 1-0.9960=0.004".

b. We should find "P\{465.1<X<1622\}."

"P\{465.1<X<1622\}=F(1622)-F(465.1)."

"F(465.1)=\Phi(\frac{465.1-703}{347})\approx \Phi(-0.686)\approx 0.2464."

"\text{So }P\{465.1<X<1622\}=F(1622)-F(465.1)\approx\\
\approx 0.9960-0.2464=0.7496."